1.
Homework Figure out the all possible offsprings from the couple which female with balanced reciprocal translocation of t (1p21;6q21) and male with normal male karyotype after reading some references.
2.
media%2F824%2F8245dbad-bdc9-49c8-a78c-95
3.
media%2F713%2F7136b880-9e8c-4f0e-8299-f9
4.
media%2F5b6%2F5b62f9cc-407c-46cb-a4c5-17
5.
media%2F7f3%2F7f3fe399-1e9b-4851-90b9-24
6.
media%2F5a5%2F5a5f49c0-cd52-44c9-910f-c8
7.
media%2F4bb%2F4bb03a2e-6762-4478-9e83-8d
Answer #1

Q1)The probability rate estimation for unfavourable pregnancies

An external file that holds a picture, illustration, etc. Object name is IJHG-18-250-g005.jpg

P= risk value

a= number of unfavorable pregnancies after ascertainment correction

s=standard deviation

n=number of all pregnancies after ascertainment correction

when n=0

the maximal risk estimate m, corresponding to the upper limit of the risk interval, is calculated using the formula:

An external file that holds a picture, illustration, etc. Object name is IJHG-18-250-g006.jpg

as reciprocal translocation is involved basic punnett square according to mendelian principles cannot be followed, was there any statistical reference ?around 2% of the population generally gets affected.

Q2Threshold trait/multifactorial/quantitative traits do not follow mendelian inheritance pattern. This traits are operated as a cumulative result of gene and environment

there are 36 genetic risk variants for predisposition of coronary artery disease (the first to be identified was 9p21. Several different genes like LDLR, APOE,PCSK9 can be involved.

Other than genes environmental factors like smoking,diet,exercise,mental stress may contribute to the underlying liability of a person to devolop such diseases.

Q3) 4 size determining genes were involved as

8/2012 = approx 1/256= 1/(4)4

Q4) broad sense heritibility= H2= Vg/Vt= 3.17/6.08=0.52

Q5) B2= R/s = (12.5-10)/ (15-10)=2.5/5=0.5

The fishermen should have the knowledge of selecting breeds with high growth increase rate.

Q6) obviously from the data it is clearly evident that alcoholism has genetics involved. Monozygotic twins almost got the half concordance of dizygotic twins. So it is clear that alcoholism is genetically inherited.

q7) The gene will be isolated , purified then its copies after amplified via PCR can be studied.